The question is: An investment will pay $100 at the end of each of the next 3 years, $200 at the end of Year 4, $300 at the end of Year 5, and $600 at the end of Year 6.

1. If other investments of equal risk earn 6% annually, what is its present value? Round your answer to the nearest cent.

2. If other investments of equal risk earn 6% annually, what is its future value? Round your answer to the nearest cent.
Answer #1

a.Present value=Cash flows*Present value of discounting factor(rate%,time period)

=100/1.06+100/1.06^2+100/1.06^3+200/1.06^4+300/1.06^5+600/1.06^6

=$1072.87(Approx).

b.We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=1072.87*(1.06)^6

=$1521.89(Approx).